2 horses are 10 km. apart and start moving towards each other, 1 horse @ 5.5km the other @ 4.5km. A horsefly moving @30km goes back and forth between them.
Question "How far will the horsefly travel when the horses meet?"

Answer by stanbon(75887)

1st find the time required for the horses to meet.
Let distance traveled by the 5.5 km horse be "x".
Then distance traveled by the 4.5 km horse will be
"10-x"
Formula: time = distance/rate
EQUATION:
Time for 5.5 horse = Time for 4.5 horse.
x/5.5 = (10-x)/4.5
4.5x = 55-5.5x
10x = 55
x = 11 km (distance for 5.5 km horse
So, time for 5.5 km horse is 11/5.5 = 2 hr.
In two hours the horsefly will go 60 km.
Cheers,
Stan H.
So